# Generative Software Engineering

### v1

@@ 2.2.2. Metric
where they generate $n\geqslant k$ samples per task, count the number of correct samples $c\leqslant n$ that pass unit tests, and calculate the unbiased estimator. Table 4 shows the pass@k metrics reported in some papers of the datasets we investigated.
Table 4. The pass@k metrics reported in some papers of the datasets we investigated
Datasets HumanEval MBPP CodeScope ClassEval
Metric Pass@1 Pass@10 Pass@100 Pass@1 Pass@5 Pass@1 Pass@3 Pass@5
GPT-3.5 48.1% - - 52.2% 22.13% 29.6% 34.9% 36.0%
GPT-4 67.0% - - - 35.29% 37.6% 41.3% 42.0%
WizardCoder - - - - 3.11% 12.2% 20.0% 23.0%
StarCoder 30.4% - - 49.0% 0.48% 10.2% 12.7% 14.0%
PaLM 26.2% - 76.2% 36.8% 2.30% - - -
Codex 28.81% 46.81% 72.31% - - - - -
Vicuna - - - - 0.36% 3.0% 3.6% 4.0%
LLAMA 2 30.5% 59.4% 87.0% 45.4% 0.72% - - -
[…]
@@ 2.2.3. Methods
Table 5 shows the performance of the LLM code generation information enhancement methods we summarized. Because different methods are based on different LLM implementations, it is meaningless to directly compare the performance of these methods. On the contrary, the performance improvement rates of their methods compared to their based LLMs are more meaningful.
Table 5. The pass@1 metrics of the method to improve LLM code generation
Type Method Dataset Based LLM Before After Improvement rate
Guiding requirement refinement ClarifyGPT HumanEval ChatGPT 64.63% 74.39% 15.10%
GPT-4 78.86% 87.80% 11.34%
MBPP ChatGPT 65.57% 74.08% 12.98%
GPT-4 70.96% 78.89% 10.89%
Self-collaboration HumanEval ChatGPT 57.3% 74.4% 29.9%
MBPP ChatGPT 52.2% 68.2% 30.7%
Guiding programming thinking. TIP HumanEval ChatGPT 52.47% 60.64% 15.57%
MBPP ChatGPT 40.00% 46.98% 17.45%
SKCODER
(pre-trained)
SKCODER SKCODER 20.0% 27.74% 38.7%
Error feedback Self-Edit HumanEval GPT-3 34.76% 39.63% 14.01%
Requirements-guided code generation. Powerful capabilities have been demonstrated by language models such as ChatGPT in automatically generating code from provided natural language requirements. However, in real-world practice, user-written requirements may be ambiguous or incomplete. Currently, LLMs struggle to handle such unclear user requirements, and generating code solutions directly from ambiguous requirements may deviate from the user’s original intent. To bridge this gap, Mu et.al. (Mu et al., 2023) introduce a method called ClarifyGPT, designed to enhance code generation by having LLM identify ambiguous requirements and propose targeted requirement clarification questions. Specifically, ClarifyGPT first inputs the requirements into LLM to generate multiple solutions. It then executes these solutions with pre-prepared test cases. If the output results of these solutions are consistent, the requirements are considered clear. Conversely, if the solutions output different results, indicating ambiguous requirements, ClarifyGPT guides LLM to generate targeted clarification questions. Upon receiving user responses, ClarifyGPT reconstructs the clarified requirements prompt and inputs it into LLM to generate code solutions.
@@
Guiding programming thinking. Programmers typically start implementing code by first writing high-level solution pseudocode, which is then refined into a concrete solution. Simultaneously, when faced with challenging code, most programmers seek similar solutions and improvements on programming forums like Stack Overflow. Inspired by this, researchers have also explored incorporating human programming thinking into Large Language Models.
@@
Domain-specific code generation. In real-world development scenarios, developers often customize APIs or classes in their projects, leading to a demand for code generation in such contexts. For instance, this includes generating code for specific library function calls or handling exceptions that may arise from specific APIs. Researchers have proposed some code generation methods under domain-specific settings.
@@
Error feedback.
Based on the experience of human developers in debug code with error reporting information, researchers fed back the errors in the code generated by LLM to LLM, so that LLM found the defects in the generated code, and modified the defect code to generate the correct code.
@@ 2.3.1. Datasets
TL-CodeSum Java A Java code summarization datasets contain 87k pairs which collected from github projects from 2015 to 2016.
@@ 2.4.2. Evaluation Criteria
Code Coverage. Code coverage measures the extent to which test cases cover the code under test and indicates how many different program paths or statements the generated test oracle or assertions cover. Higher coverage generally indicates more comprehensive testing. It includes metrics such as statement coverage, branch coverage, and condition coverage. Higher code coverage indicates that the test cases cover a significant portion of the code. However, high coverage doesn’t guarantee test effectiveness or completeness.
Path Coverage. Path coverage evaluates whether test cases cover all possible execution paths within a software program. High path coverage enhances the thoroughness of test cases, but it may not cover all complex paths.
Fault Detection Rate. Fault detection rate measures the success of test cases in identifying defects or issues within the software system. A high fault detection rate suggests that the test cases are effective in detecting potential problems, though it doesn’t guarantee to find all issues.
Retesting Rate. Retesting rate indicates whether running the same test cases multiple times produces consistent results. A low retesting rate implies good repeatability of test cases and stable test results.
Test Case Execution Time. Test case execution time is the time taken for a test suite to complete execution. Assessment: Shorter test case execution times may indicate efficient testing, but it’s crucial to balance speed with the quality of test cases.
Accuracy. A measure of how well the generated test oracle or assertions match the actual expected results. High accuracy means that the generated content is more consistent with the actual desired results.
2.4.3. Method: Comprehensive Test Case Generation
Methods for pre-training and fine-tuning training based on LLM. Tufano et al. (Tufano et al., 2020) proposed ATHENATEST (based on the BART transformer architecture), which aims to generate unit test cases by learning real-world focal-methods and test cases written by developers. They formulate unit test case generation as a sequence-to-sequence learning task, employing a two-step training process that includes denoising pre-training on a large unsupervised Java corpus and supervised fine-tuning of the downstream task of generating unit tests. Li et al. (Li et al., 2022) proposed a large-scale code pre-training language model AlphaCode (encoder-decoder architecture based on transformer), which was pre-trained on a large-scale code corpus collected by GitHub. In the test case generation task, they fine-tune AlphaCode on the corresponding dataset.
@@ 2.4.3. Method: Comprehensive Test Case Generation
Methods based on LLM and prompt engineering.
Patrick et al. (Bareiß et al., 2022) used a large-scale code pre-training language model CodeX and few-shot learning for generating test cases. Their method is mainly divided into three steps: (1)instance extraction (i.e. some examples of test case generation are added to the prompt (i.e. few-shot learning)). (2) enter the designed prompt into the Codex. (3) post-process generated test cases, such as de-duplication and discarding test cases that do not pass compilation.
@@
Methods based on LLM, prompt engineering and strategy.
Schäfer et al. (Schäfer et al., 2023) proposed TESTPILOT, an adaptive JavaScript test generation tool based on OpenAI’s gpt3.5-turbo LLM, which can automatically generate unit tests for methods in a given project’s API. TESTPILOT embeds contextual information about the function under test into the prompt, specifically its signature, additional documentation comments (if any), any usage examples in the project documentation, and the function’s source code. Additionally, if a generated test fails, TESTPILOT will adaptively create a new prompt embedding the test and failure messages to guide the model in fixing the problematic test.
@@
2.4.4. Method: Test Oracle and Assert Statements Generation in Test Case
Methods for fine-tuning training based on LLM. White et al. (White and Krinke, 2020) proposed a method REASSERT to automatically generate JUnit test assertions based on the pre-trained model Reformer. REASSERT first collects test and tested method pairs from the target project via test-to-code traceability links. For each test and tested method pair, we extract assertion statements from the test method and concatenate them to generate a string of assertion statements associated with the method under test. The method and assertion strings under test are then processed into sequences of input and output tokens, called method sequences and assertion sequences respectively. These sequences are used to train the pre-trained model Reformer (a transformer-based model). The resulting sequence of assertions is processed into syntactically correct code that can be inserted directly into the method’s test case.
@@ 2.4.4. Method: Test Oracle and Assert Statements Generation in Test Case
Methods for pre-training and fine-tuning training based on LLM. Tufano et al. (Tufano et al., 2022) proposed a method to generate accurate and useful assertion statements based on the BART transformer model to support developers in writing unit test cases. This method uses two pre-training stages: (1) pre-training: semi-supervised pre-training on a large corpus of English text and code pre-training on the Java source code dataset. (2) fine-tuning on the test cases and the labeled dataset. Mastropaolo et al. (Mastropaolo et al., 2021) explored the performance of pre-training and fine-tuning for supporting code-related tasks based on T5 LLM. They pre-trained on the CodeSearchNet dataset, where an abstract version of each mined Java method was created using the src2abs tool. For the assertion statement generation task of test cases, they used the open source dataset (Watson et al., 2020) for fine-tuning training.
@@ 2.7. Code Translation
One-to-one code translation CodeTrans (Lu et al., 2021) Java/C#
@@ 2.7.2. Evaluation Criteria
Error
Description
Translation Error (TE)
Errors in translating code statements, including mistakes in assignment statements, conditional statements, loop conditions, array lookups, spaces, or other logical statements.
Language Error (LE)
included Java code snippets within Python or failed to appropriately translate Java language idioms into Pythonic expressions.
Spurious Error (SE)
Participant introduced functionality that was not part of the original Java program, such as by defining new methods.
Code Omission Error (COE)
neglected to translate a method or code statements within a method, providing a trivial implementation (e.g., using pass, return None, print(’not implemented’), etc.).
Documentation Omission Error (DOE)
failed to provide translation for a function’s documentation (e.g., Javadoc comment).
Correctness Error (CE)
translation of a method was inaccurate (e.g., did not pass unit tests).
@@ 3. Challenge
To overcome these challenges and further advance research on LLMs in the code generation domain, here are some future research directions:
@@
Code Generation: using pre-trained models or LLMs for code generation is a field with potential and challenges. Existing techniques can be categorized into (1) guided approaches: using various techniques and methods to guide the learning and generation process of LLMs, improving the accuracy, performance, and quality of code generation, and (2) error-feedback approaches: providing error feedback to LLMs to help them discover and improve defects in the generated code. This includes self-editing and self-collaboration methods, using execution results and test reports to guide LLMs in code debugging and improvement. We summarize some challenges faced by these techniques, including but not limited to (1) Data and annotation challenges: To train and evaluate the code generation capability of LLMs, existing evaluation datasets require manually constructed test cases, which require a large amount of high-quality data and corresponding annotations. Acquiring and annotating such data can be time-consuming and challenging tasks. (2) Lack of domain knowledge: Although LLMs perform well in language understanding and generation, they may lack in-depth knowledge of specific domains, such as the usage of specific APIs and functions, handling specific exceptions, etc. For domain-specific code generation tasks, LLMs may require additional training and guidance to ensure that the generated code complies with domain best practices and constraints. (2) Assurance of code quality and generation robustness: The generated code may have quality and performance issues, such as vulnerable code. Improving the quality of generated code and ensuring generation robustness are important challenges.
To overcome these challenges and further advance research in LLMs for code generation, here are some feasible future research directions: (1) Data augmentation and automatic annotation: Develop more effective methods to acquire and annotate large-scale code generation data. This can include rule-based data augmentation techniques, code transformations, and automatic annotation methods to alleviate the burden of data acquisition and annotation. (2) Introduction of domain knowledge: Explore how to incorporate domain-specific knowledge into the training and generation process of LLMs to improve code quality and compliance with domain requirements. This can include guidance from domain experts, domain-specific training data, and the construction of knowledge bases. (3) Quality and performance assurance: Develop techniques and methods to ensure that the generated code has high quality and good performance. This can involve the application of code analysis and verification techniques, as well as automatic correction and rewriting based on constraints and best practices. These research directions contribute to overcoming the current challenges in LLM code generation and further advancing the field.
Code summarization: Code summarization models undergo training and evaluation using extensive datasets. However, the practical applicability of these models to assist developers in generating summaries for their specific projects remains unclear. In essence, there is a need for code summarization models to exhibit greater adaptability to the unique characteristics of individual projects within different domains. This suggests that there is ample room for exploration and improvement in tailoring code summarization models to cater to the diverse needs of developers working on varied projects.
Although we can establish numerous filtering rules, human-written comments exhibit considerable variability, introducing bias to code summarization models. Large language models can leverage elements such as code structure, function names, and variable names to infer the likely functionality of the code and generate concise summaries.(Sun et al., 2023b) Can summaries generated by such large language models be considered as supplementary information or even directly utilized as ground truth by humans? There is still much to explore in this regard.

### v2

@@ 2.2.2. Metric
where they generate $n\geqslant k$ samples per task, count the number of correct samples $c\leqslant n$ that pass unit tests, and calculate the unbiased estimator.
RPs@k, RDs@k, and RRs@k: To assess the robustness of generated code, Wang et.al. (Wang et al., 2023d) defined robustness metrics for code generation models in ReCode: Robust Passs@k (RPs@k), Robust Drops@k (RDs@k), and Robust Relatives@k (RRs@k). The calculation methods of these three metrics are shown as formulas 2 to 4:
@@ 2.2.3. Methods
Requirements-guided code generation: Powerful capabilities have been demonstrated by language models such as ChatGPT in automatically generating code from provided natural language requirements. However, in real-world practice, user-written requirements may be ambiguous or incomplete. Currently, LLMs struggle to handle such unclear user requirements, and generating code solutions directly from ambiguous requirements may deviate from the user’s original intent. To bridge this gap, Mu et.al. (Mu et al., 2023) introduce a method called ClarifyGPT, designed to enhance code generation by having LLM identify ambiguous requirements and propose targeted requirement clarification questions. Specifically, ClarifyGPT first inputs the requirements into LLM to generate multiple solutions. It then executes these solutions with pre-prepared test cases. If the output results of these solutions are consistent, the requirements are considered clear. Conversely, if the solutions output different results, indicating ambiguous requirements, ClarifyGPT guides LLM to generate targeted clarification questions. Upon receiving user responses, ClarifyGPT reconstructs the clarified requirements prompt and inputs it into LLM to generate code solutions.
@@
Guiding programming thinking: Programmers typically start implementing code by first writing high-level solution pseudocode, which is then refined into a concrete solution. Simultaneously, when faced with challenging code, most programmers seek similar solutions and improvements on programming forums like Stack Overflow. Inspired by this, researchers have also explored incorporating human programming thinking into Large Language Models.
@@
Domain-specific code generation: In real-world development scenarios, developers often customize APIs or classes in their projects, leading to a demand for code generation in such contexts. For instance, this includes generating code for specific library function calls or handling exceptions that may arise from specific APIs. Researchers have proposed some code generation methods under domain-specific settings.
@@
Error feedback: Based on the experience of human developers in debug code with error reporting information, researchers fed back the errors in the code generated by LLM to LLM, so that LLM found the defects in the generated code, and modified the defect code to generate the correct code.
@@ 2.3.1. Datasets
TL-CodeSum Java
A Java code summarization datasets contain 87k pairs
which collected from github projects from 2015 to 2016.
@@ 2.4.2. Evaluation Criteria
Code Coverage: Code coverage measures the extent to which test cases cover the code under test and indicates how many different program paths or statements the generated test oracle or assertions cover. Higher coverage generally indicates more comprehensive testing. It includes metrics such as statement coverage, branch coverage, and condition coverage. Higher code coverage indicates that the test cases cover a significant portion of the code. However, high coverage doesn’t guarantee test effectiveness or completeness.
Path Coverage: Path coverage evaluates whether test cases cover all possible execution paths within a software program. High path coverage enhances the thoroughness of test cases, but it may not cover all complex paths.
Fault Detection Rate: Fault detection rate measures the success of test cases in identifying defects or issues within the software system. A high fault detection rate suggests that the test cases are effective in detecting potential problems, though it doesn’t guarantee to find all issues.
Retesting Rate: Retesting rate indicates whether running the same test cases multiple times produces consistent results. A low retesting rate implies good repeatability of test cases and stable test results.
Test Case Execution Time: Test case execution time is the time taken for a test suite to complete execution. Assessment: Shorter test case execution times may indicate efficient testing, but it’s crucial to balance speed with the quality of test cases.
Accuracy: A measure of how well the generated test oracle or assertions match the actual expected results. High accuracy means that the generated content is more consistent with the actual desired results.
2.4.3. Method.
Many test case generation methods have been proposed in recently years.
Methods for pre-training and fine-tuning training based on LLM: Tufano et al. (Tufano et al., 2020) proposed ATHENATEST (based on the BART transformer architecture), which aims to generate unit test cases by learning real-world focal-methods and test cases written by developers. They formulate unit test case generation as a sequence-to-sequence learning task, employing a two-step training process that includes denoising pre-training on a large unsupervised Java corpus and supervised fine-tuning of the downstream task of generating unit tests. Li et al. (Li et al., 2022) proposed a large-scale code pre-training language model AlphaCode (encoder-decoder architecture based on transformer), which was pre-trained on a large-scale code corpus collected by GitHub. In the test case generation task, they fine-tune AlphaCode on the corresponding dataset.
@@
Methods based on LLM and prompt engineering: Patrick et al. (Bareiß et al., 2022) used a large-scale code pre-training language model CodeX and few-shot learning for generating test cases. Their method is mainly divided into three steps: (1)instance extraction (i.e. some examples of test case generation are added to the prompt (i.e. few-shot learning)). (2) enter the designed prompt into the Codex. (3) post-process generated test cases, such as de-duplication and discarding test cases that do not pass compilation.
@@
Methods based on LLM, prompt engineering and strategy: Schäfer et al. (Schäfer et al., 2023) proposed TESTPILOT, an adaptive JavaScript test generation tool based on OpenAI’s gpt3.5-turbo LLM, which can automatically generate unit tests for methods in a given project’s API. TESTPILOT embeds contextual information about the function under test into the prompt, specifically its signature, additional documentation comments (if any), any usage examples in the project documentation, and the function’s source code. Additionally, if a generated test fails, TESTPILOT will adaptively create a new prompt embedding the test and failure messages to guide the model in fixing the problematic test.
@@
2.4.4. Method:
There are many methods regarding the test oracle and assert statements generation in test case.
White et al. (White and Krinke, 2020) proposed a method REASSERT to automatically generate JUnit test assertions based on the pre-trained model Reformer. REASSERT first collects test and tested method pairs from the target project via test-to-code traceability links. For each test and tested method pair, we extract assertion statements from the test method and concatenate them to generate a string of assertion statements associated with the method under test. The method and assertion strings under test are then processed into sequences of input and output tokens, called method sequences and assertion sequences respectively. These sequences are used to train the pre-trained model Reformer (a transformer-based model). The resulting sequence of assertions is processed into syntactically correct code that can be inserted directly into the method’s test case.
@@ 2.4.4. Method:
Tufano et al. (Tufano et al., 2022) proposed a method to generate accurate and useful assertion statements based on the BART transformer model to support developers in writing unit test cases. This method uses two pre-training stages: (1) pre-training: semi-supervised pre-training on a large corpus of English text and code pre-training on the Java source code dataset. (2) fine-tuning on the test cases and the labeled dataset. Mastropaolo et al. (Mastropaolo et al., 2021) explored the performance of pre-training and fine-tuning for supporting code-related tasks based on T5 LLM. They pre-trained on the CodeSearchNet dataset, where an abstract version of each mined Java method was created using the src2abs tool. For the assertion statement generation task of test cases, they used the open source dataset (Watson et al., 2020) for fine-tuning training.
@@ 2.5.2. Evaluation Criteria
When verifying the effectiveness of patch generation methods, qualitative or quantitative indicators are often used to measure them. Evaluating the effectiveness of code patch generation techniques typically involves multiple metrics, and the following are some commonly used evaluation metrics:
Repair accuracy: Repair accuracy refers to the proportion of the number of bugs that the model can correctly fix in the provided defective code test set to the total number of defects in the test set. The calculation method is as follows:
(5) RA=Number of Successfully Repaired DefectsTotal Number of Attempted $\text{RA}=\frac{\text{Number of Successfully Repaired Defects}}{\text{Total Number of Attempted Defects}}\times 100\%$
Where RA is Repair Accuracy. "successfully repaired defects" means that for a defect test suite where each defect may have multiple test cases, only when the generated patch can pass all test cases without introducing new defects, can it be considered as successfully fixing this defect. This usually requires manual verification or additional verification steps to determine. The repair accuracy directly reflects the quality of the generated patches, and high accuracy means that the generated patches are more reliable in repairing defects.
Overlap analysis: Overlap analysis is a metric used in the field of APR to evaluate the diversity and originality of generated patches. During the process of APR, a system may generate multiple patches that fix the same defect. These patches may have overlapping fix strategies, i.e. they may modify the same lines of code or adopt similar fix patterns. Researchers often use this metric to analyze how well their proposed patch generation methods can complement existing methods. If a patch generated by one method can fix more defects that cannot be fixed by other methods, it indicates that it has more general repair capabilities. Overlap analysis is calculated by counting the number of bugs in the overlapping parts of bugs fixed by different technologies and displaying them visually using Venn diagrams.
Compilation rate: During the automated program repair process, a large number of candidate patches may be generated. Not all generated patches can pass compilation. Only patches that can be successfully compiled are likely to be effective fixes, and only successfully compiled patches need to undergo subsequent testing (for example, running test cases to check whether the defect is fixed). The patch compilation rate is a metric used to measure the extent to which the generated patches can be successfully compiled by the compiler. The compilation rate directly affects whether the patch can be further tested and deployed, so it is an important indicator to measure the quality of the patch. The compilation rate of a patch is usually calculated according to the following formula:
(6) CR=Number of Patches that Compile SuccessfullyTotal Number of Generated $\text{CR}=\frac{\text{Number of Patches that Compile Successfully}}{\text{Total Number of Generated Patches}}\times 100\%$
Where CR is the Compilation rate.
@@ 2.7. Code Translation
One-to-one
code translation
CodeTrans (Lu et al., 2021) Java/C#
@@ 2.7.2. Evaluation Criteria
Error Description
Translation Error (TE) Errors in translating code statements, including mistakes in assignment statements, conditional statements, loop conditions, array lookups, spaces, or other logical statements.
Language Error (LE) included Java code snippets within Python or failed to appropriately translate Java language idioms into Pythonic expressions.
Spurious Error (SE) Participant introduced functionality that was not part of the original Java program, such as by defining new methods.
Code Omission Error (COE) neglected to translate a method or code statements within a method, providing a trivial implementation (e.g., using pass, return None, print(’not implemented’), etc.).
Documentation Omission Error (DOE) failed to provide translation for a function’s documentation (e.g., Javadoc comment).
Correctness Error (CE) translation of a method was inaccurate (e.g., did not pass unit tests).
@@ 3. Challenge
To overcome these challenges and further advance research on LLMs in the requirement generation domain, here are some future research directions:
@@
Code Generation: using pre-trained models or LLMs for code generation is a field with potential and challenges. Existing techniques can be categorized into (1) guided approaches: using various techniques and methods to guide the learning and generation process of LLMs, improving the accuracy, performance, and quality of code generation, and (2) error-feedback approaches: providing error feedback to LLMs to help them discover and improve defects in the generated code. This includes self-editing and self-collaboration methods, using execution results and test reports to guide LLMs in code debugging and improvement. We summarize some challenges faced by these techniques, including but not limited to
  1. 1)
    Data and annotation challenges: To train and evaluate the code generation capability of LLMs, existing evaluation datasets require manually constructed test cases, which require a large amount of high-quality data and corresponding annotations. Acquiring and annotating such data can be time-consuming and challenging tasks. Lack of domain knowledge: Although LLMs perform well in language understanding and generation, they may lack in-depth knowledge of specific domains, such as the usage of specific APIs and functions, handling specific exceptions, etc. For domain-specific code generation tasks, LLMs may require additional training and guidance to ensure that the generated code complies with domain best practices and constraints.
  2. 2)
    Lack of domain knowledge: Although LLMs perform well in language understanding and generation, they may lack in-depth knowledge of specific domains, such as the usage of specific APIs and functions, handling specific exceptions, etc. For domain-specific code generation tasks, LLMs may require additional training and guidance to ensure that the generated code complies with domain best practices and constraints.
  3. 3)
    Assurance of code quality and generation robustness: The generated code may have quality and performance issues, such as vulnerable code. Improving the quality of generated code and ensuring generation robustness are important challenges.
To overcome these challenges and further advance research in LLMs for code generation, here are some feasible future research directions:
  1. 1)
    Data augmentation and automatic annotation: Develop more effective methods to acquire and annotate large-scale code generation data. This can include rule-based data augmentation techniques, code transformations, and automatic annotation methods to alleviate the burden of data acquisition and annotation.
  2. 2)
    Introduction of domain knowledge: Explore how to incorporate domain-specific knowledge into the training and generation process of LLMs to improve code quality and compliance with domain requirements. This can include guidance from domain experts, domain-specific training data, and the construction of knowledge bases.
  3. 3)
    Quality and performance assurance: Develop techniques and methods to ensure that the generated code has high quality and good performance. This can involve the application of code analysis and verification techniques, as well as automatic correction and rewriting based on constraints and best practices.
These research directions contribute to overcoming the current challenges in LLM code generation and further advancing the field.
3.0.1. Challenge and Future
Challenge
Language Specificity:
Each programming language has its own syntax, semantics, and coding conventions. Models trained on datasets for one language may not generalize well to others due to differences in language structure and usage patterns. Thus, separate datasets are required for each language, increasing the demand for diverse datasets across multiple programming languages.
Project Specificity: Software projects are project-specific and may involve specialized terminology, libraries, or frameworks. Generating accurate summaries for such code requires models that can understand and incorporate project-specific knowledge. Adapting generic code summarization techniques to unique project while ensuring relevance and accuracy is a non-trivial task.
Evaluation Metrics: Assessing the quality of generated code summaries remains an open challenge. Traditional metrics like BLEU, ROUGE and METEOR used in natural language generation tasks may not fully capture the effectiveness of code summaries in aiding developer understanding and productivity. Developing robust evaluation metrics that account for the unique characteristics of code summarization is essential for advancing the field.
Future Customized Summarization: Since different developers may have varying preferences, expertise levels, and information needs, personalized code summarization techniques can help tailor summaries to individual users. Future research could explore methods for incorporating user feedback, preferences, and historical interactions to personalize the summarization process.
Multi-Granularity Summarization: Codebases consist of code at various levels of granularity, including functions, classes, modules, and entire projects. Generating summaries that are informative and coherent across different levels of granularity poses a challenge. Future research could investigate techniques for multi-granularity summarization, where summaries are generated at different levels of abstraction and detail to cater to different information needs and use cases. This could involve hierarchical summarization approaches, abstraction mechanisms, and content selection strategies tailored to different levels of code granularity.
Explainable Summarization: Investigating methods for generating explainable summaries that not only describe what the code does but also provide insights into why certain decisions were made during its implementation. This involves providing explanations or rationales alongside code summaries to aid developer understanding.